# The pattern block method for generating random numbers : Reformulation and generalization of the Ziggurat method using conditional random variables

### 17 Mar 2026 v1

@@ 1 Introduction
Random number generation is a foundational component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness through numerical examples. However, the random variable underlying the Ziggurat algorithm has not been formally defined. Therefore, in this study, we reformulate and generalize the Ziggurat method using conditional random variables. Conditional random variables have played an important role in [7, 8, 9, 10].
@@ 2 The pattern block method
Using (5), (14) and Fubini’s theorem, we obtain (12). ∎
@@ 3 Applications to generate 11-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})=(\mathbb{R},\mathcal{B}(\mathbb{R}))$ holds and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then $\nu$ defined by (2) is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
@@ 4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is given by $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$
@@
Fig. 9 shows the density scale histogram of $500,000$ random vectors and the density function $y=f(x_{1},x_{2})$.
@@ 5 Conclusion
We introduced the pattern block method, which is a generalization of the Ziggurat method. In our future work, we aim to develop an algorithm that identifies pattern block with high adoption rate and low search cost.

### 19 Mar 2026 v2

@@ 1 Introduction
Random number generation is a foundational component of Monte Carlo methods. Recently, [1, 2, 3, 4, 5, 6] developed the Ziggurat method for generating 11-dimensional normally distributed random numbers and demonstrated its effectiveness through numerical examples. However, the random variable underlying the Ziggurat algorithm has not been defined. Therefore, in this study, we reformulate and generalize the Ziggurat method using conditional random variables. Conditional random variables have played an important role in [7, 8, 9, 10].
@@ 2 The pattern block method
Using (5), (14) and the Fubini’s theorem, we obtain (12). ∎
@@ 3 Applications to generate 11-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is $(\mathbb{R},\mathcal{B}(\mathbb{R}))$ and $\mu$ is the 11-dimensional Lebesgue measure on $(\mathbb{R},\mathcal{B}(\mathbb{R}))$. Then $\nu$ defined by (2) is the 22-dimensional Lebesgue measure on $(\mathbb{R}^{2},\mathcal{B}(\mathbb{R}^{2}))$. For every $C\in\mathcal{B}(\mathbb{R}^{2})$, $|C|$ denotes the 22-dimensional Lebesgue measure $\nu(C)$.
@@ 4 Application to generate 22-dimensional random numbers
In this section, we assume that $(E,\mathcal{E})$ is given by $E=[\underline{a},\overline{a}]\times[\underline{a},\overline{a}]$ and $\mathcal{E}=\mathcal{B}(E)$, where $\underline{a}=-4$ and $\overline{a}=4$. Let $\mu$ be the 22-dimensional Lebesgue measure on $(E,\mathcal{E})$. Then $\nu$ is the 33-dimensional Lebesgue measure on $(E\times\mathbb{R},\mathcal{B}(E\times\mathbb{R}))$. For every $C\in\mathcal{B}(E\times\mathbb{R})$, $|C|$ denotes the 33-dimensional Lebesgue measure $\nu(C)$. Let $f(x_{1},x_{2})$ be the density function on $(E,\mathcal{E})$
@@
Fig. 9 and Fig. 9 show the density scale histogram of $500,000$ random vectors and the density function $y=f(x_{1},x_{2})$, respectively.
@@ 5 Conclusion
We introduced the pattern block method, which is a generalization of the Ziggurat method. In our future work, we aim to develop an algorithm that identifies pattern block of a given density function with high adoption rate and low search cost.